A 1.0-kg bar of copper is heated at atmospheric pressure so that its temperature increases from 20°C to 50°C. How much energy is transferred to the copper bar by heat?


1
Expert's answer
2021-05-17T14:43:01-0400

Specific heat of copper C=387J/kg degreeCm=1kg

Δq=mcΔT=mc(T2-T1)=1*387*(50-30)=11610J